How do I convert radians to degrees? Is it pi/180 or 180/pi? I get them confused! Thanks
Found 2 solutions by stanbon, tinbar:
Answer by stanbon(75887)

How do I convert radians to degrees? Is it pi/180 or 180/pi?
---
This is how I remember:
If you multiply by 180 on the top, you end up with degrees.
---
If you multiply by pi on the top, you end up with radians.
---
Example:
(36 degrees)*(pi/180) = 36pi/180 = (1/5)pi radians
====
(14 radians)*(180/pi) = (14*180/pi) = 802.14 degrees
=========================================================
Cheers,
Stan H.

Answer by tinbar(133)

the best way to remember this kind of stuff is to consider units. Radians are measured in "pi" units so to speak, so if it were times 180/pi it wouldn't really make sense since you are getting something/pi units, however, if you multiply it by pi/180 then you get some pi units.
You can also consider specific angles you may know in both degrees and radians and see which operation makes sense. For instance, I know that 360 degrees is 2*pi. If I perform the operation 360*(pi/180) I get 2*pi, which suggests that multiplying degrees by pi/180 converts it to radians